John's and Mary's ages sum up to 25.Eight years ago,John was twice as old as Mary.How old are they now?
Answer by macston(5194)

.
J=John's age; M=Mary's age
.
J+M=25
J=25-M
.
J-8=2(M-8)
(25-M)-8=2M-16
17-M=2M-16
33-M=2M
33=3M
11=M
.
Mary is 11 years old.
.
J=25-M=25-11=14
.
John is 14 years old.
.
ANSWER: John is 14 years old and Mary is 11 years old
.
CHECK:
J-8=14-8=6
8 years ago, John was 6 years old.
.
M-8=11-8=3
8 years ago, Mary was 2 years old.
.
8 years ago, John was twice as old as Mary:
6yrs=2(3yrs)
6yrs=6yrs
.